# Determination of Magnetic Symmetries by Convergent Beam Electron Diffraction

### v2

@@ Appendix C Mappings of magnetic point groups to magnetic CBED groups
$32$ $32^{\prime}$ $3m^{\prime}_{R}$ $2^{\prime}$ $m^{\prime}_{R}$ $1$
$3m$ ($C_{3v}$) $3m$ $3m$ $1_{R}$ $m$ $1$
$3m$ $3m1^{\prime}$ $3m1^{\prime}$ $1_{R}1^{\prime}$ $m1^{\prime}$ $11^{\prime}$
$3m$ $3m^{\prime}$ $3m^{\prime}$ $1^{\prime}_{R}$ $m^{\prime}$ $1$
@@
[…]
$6$ ($C_{6}$) $6$ $6$ $m_{R}$ $1$
$6$ $61^{\prime}$ $61^{\prime}$ $m_{R}1^{\prime}$ $11^{\prime}$
$6$ $6^{\prime}$ $6^{\prime}$ $m^{\prime}_{R}$ $1$
$\bar{6}$ ($C_{3h}$) $\bar{6}$ $31_{R}$ $m$ $1$
@@
$\bar{6}$ $\bar{6}^{\prime}$ $31^{\prime}_{R}$ $m^{\prime}$ $1$
@@
$622$ $6^{\prime}22^{\prime}$ $6^{\prime}m_{R}m^{\prime}_{R}$ $2m^{\prime}_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m^{\prime}_{R}$ $m^{\prime}_{R}$ $m_{R}$ $1$
$622$ $62^{\prime}2^{\prime}$ $6m^{\prime}_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m_{R}$ $m^{\prime}_{R}$ $m^{\prime}_{R}$ $1$
$6mm$ ($C_{6v}$) $6mm$ $6mm$ $m1_{R}$ $m1_{R}$ $m_{R}$ $m$ $m$ $1$
$6mm$ $6mm1^{\prime}$ $6mm1^{\prime}$ $m1_{R}1^{\prime}$ $m1_{R}1^{\prime}$ $m_{R}1^{\prime}$ $m1^{\prime}$ $m1^{\prime}$ $11^{\prime}$
$6mm$ $6^{\prime}mm^{\prime}$ $6^{\prime}mm^{\prime}$ $m^{\prime}1_{R}$ $m1^{\prime}_{R}$ $m^{\prime}_{R}$ $m^{\prime}$ $m$ $1$
$6mm$ $6m^{\prime}m^{\prime}$ $6m^{\prime}m^{\prime}$ $m^{\prime}1^{\prime}_{R}$ $m^{\prime}1^{\prime}_{R}$ $m_{R}$ $m^{\prime}$ $m^{\prime}$ $1$
$\bar{6}m2$ ($D_{3h}$) $\bar{6}m2$ $3m1_{R}$ $m1_{R}$ $2mm$ $m$ $m_{R}$ $m$ $1$
@@
$\bar{6}m2$ $\bar{6}^{\prime}m^{\prime}2$ $3m^{\prime}1^{\prime}_{R}$ $m^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}$ $m^{\prime}$ $m_{R}$ $m^{\prime}$ $1$
$\bar{6}m2$ $\bar{6}^{\prime}m2^{\prime}$ $3m1^{\prime}_{R}$ $m^{\prime}1_{R}$ $2^{\prime}mm^{\prime}$ $m^{\prime}$ $m^{\prime}_{R}$ $m$ $1$
$\bar{6}m2$ $\bar{6}m^{\prime}2^{\prime}$ $3m^{\prime}1_{R}$ $m1^{\prime}_{R}$ $2^{\prime}mm^{\prime}$ $m$ $m^{\prime}_{R}$ $m^{\prime}$ $1$
@@
$432$ $4^{\prime}32^{\prime}$ $4^{\prime}m_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $3m^{\prime}_{R}$ $m_{R}$ $m^{\prime}_{R}$ $1$
$\bar{4}3m$ ($T_{d}$) $\bar{4}3m$ $4_{R}mm_{R}$ $m1_{R}$ $3m$ $m_{R}$ $m$ $1$
@@
$\bar{4}3m$ $\bar{4}^{\prime}3m^{\prime}$ $4^{\prime}_{R}m^{\prime}m_{R}$ $m^{\prime}1^{\prime}_{R}$ $3m^{\prime}$ $m_{R}$ $m^{\prime}$ $1$
@@ Appendix D Thermal Diffuse Scattering
This appendix details the scattering simulations carried out to assess the impact of thermal diffuse scattering (phonon scattering) on magnetic CBED. We have chosen parallel illumination to reduce the computational effort (smaller supercell) considering that the TDS background intensity is similar to the one obtained for converged illumination (CBED), see subsection V.2. We performed electron diffraction simulations for NiO including thermal diffuse scattering (TDS) using a frozen-phonon approximation [64]. Snapshots of the vibrating structure were calculated using molecular dynamics, performed using the atomic simulation environment ([65]) within Langevin dynamics. The correlated motion of atoms was described using the Orb foundational model as the interatomic potential [66]. The assumed orientation of the sample is $\left[11\bar{2}\right]_{c}$, the acceleration voltage is $300\,\mathrm{kV}$, and the size of the supercell is $29.0292\,\text{\AA}\times 29.6278\,\text{\AA}\times 1000.678\,\text{\AA}$. The thickness of the sample is the same as in the case of relevant magnetic CBED simulations. The average value of the normalized TDS intensity at the location of magnetic reflections (red dashed circles) is $2.3\times 10^{-3}$, see Fig. 8.

### v3

@@ 33 $^{\prime} 3131^{\prime}$ $11^{\prime}$
$32$ $32^{\prime}$ $3m^{\prime}_{R}$ $2^{\prime}$ $m^{\prime}_{R}$ 11
$3m$ ($C_{3v}$) $3m$ $3m$ $1_{R}$ $m$ 11
$3m$ $^{\prime} 3m13m1^{\prime}$ $1_{R}1^{\prime}$ $m1^{\prime}$ $11^{\prime}$
$3m$ $^{\prime} 3m3m^{\prime}$ $1^{\prime}_{R}$ $m^{\prime}$ 11
@@
$[0001]$ $[uv.0]$ $[uv.w]$
66 ($C_{6}$) 66 66 $m_{R}$ 11
66 $^{\prime} 6161^{\prime}$ $m_{R}1^{\prime}$ $11^{\prime}$
66 $^{\prime} 66^{\prime}$ $m^{\prime}_{R}$ 11
$\bar{6}$ ($C_{3h}$) $\bar{6}$ $31_{R}$ $m$ 11
@@ 66 $^{\prime} 66^{\prime}$ $m^{\prime}_{R}$ 11
$\bar{6}$ $\bar{6}^{\prime}$ $31^{\prime}_{R}$ $m^{\prime}$ 11
@@
$622$ $6^{\prime}22^{\prime}$ $6^{\prime}m_{R}m^{\prime}_{R}$ $2m^{\prime}_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m^{\prime}_{R}$ $m^{\prime}_{R}$ $m_{R}$ 11
$622$ $62^{\prime}2^{\prime}$ $6m^{\prime}_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $m_{R}$ $m^{\prime}_{R}$ $m^{\prime}_{R}$ 11
$6mm$ ($C_{6v}$) $6mm$ $6mm$ $m1_{R}$ $m1_{R}$ $_{R} mm$ $m$ 11
$6mm$ $^{\prime} 6mm16mm1^{\prime}$ $m1_{R}1^{\prime}$ $m1_{R}1^{\prime}$ $m_{R}1^{\prime}$ $^{\prime} m1m1^{\prime}$ $11^{\prime}$
$6mm$ $6^{\prime}mm^{\prime}$ $6^{\prime}mm^{\prime}$ $m^{\prime}1_{R}$ $m1^{\prime}_{R}$ $^{\prime}_{R} mm^{\prime}$ $m$ 11
$6mm$ $^{\prime}m^{\prime} 6mm6$$^{\prime}m^{\prime}$ $m^{\prime}1^{\prime}_{R}$ $m^{\prime}1^{\prime}_{R}$ $_{R} mm^{\prime}$ $m^{\prime}$ 11
$\bar{6}m2$ ($D_{3h}$) $\bar{6}m2$ $3m1_{R}$ $m1_{R}$ $2mm$ $m$ $_{R} mm$ 11
@@
$\bar{6}m2$ $\bar{6}^{\prime}m^{\prime}2$ $3m^{\prime}1^{\prime}_{R}$ $m^{\prime}1^{\prime}_{R}$ $2m^{\prime}m^{\prime}$ $m^{\prime}$ $_{R} mm^{\prime}$ 11
$\bar{6}m2$ $\bar{6}^{\prime}m2^{\prime}$ $3m1^{\prime}_{R}$ $m^{\prime}1_{R}$ $2^{\prime}mm^{\prime}$ $m^{\prime}$ $^{\prime}_{R} mm$ 11
$\bar{6}m2$ $\bar{6}m^{\prime}2^{\prime}$ $3m^{\prime}1_{R}$ $m1^{\prime}_{R}$ $2^{\prime}mm^{\prime}$ $m$ $^{\prime}_{R} mm^{\prime}$ 11
@@
$432$ $4^{\prime}32^{\prime}$ $4^{\prime}m_{R}m^{\prime}_{R}$ $2^{\prime}m_{R}m^{\prime}_{R}$ $3m^{\prime}_{R}$ $m_{R}$ $m^{\prime}_{R}$ 11
$\bar{4}3m$ ($T_{d}$) $\bar{4}3m$ $4_{R}mm_{R}$ $m1_{R}$ $3m$ $_{R} mm$ 11
@@
$\bar{4}3m$ $\bar{4}^{\prime}3m^{\prime}$ $4^{\prime}_{R}m^{\prime}m_{R}$ $m^{\prime}1^{\prime}_{R}$ $3m^{\prime}$ $_{R} mm^{\prime}$ 11
@@ Appendix D Thermal Diffuse Scattering
This appendix details the scattering simulations carried out to assess the impact of thermal diffuse scattering (phonon scattering) on magnetic CBED. We have chosen parallel illumination to reduce the computational effort (smaller supercell) considering that the TDS background intensity is similar to the one obtained for converged illumination (CBED), see subsection V.2. We performed electron diffraction simulations for NiO including thermal diffuse scattering (TDS) using a frozen-phonon approximation [Loane1991]. Snapshots of the vibrating structure were calculated using molecular dynamics, performed using the atomic simulation environment ([ase-paper]) within Langevin dynamics. The correlated motion of atoms was described using the Orb foundational model as the interatomic potential [ORB]. The assumed orientation of the sample is $\left[11\bar{2}\right]_{c}$, the acceleration voltage is $300\,\mathrm{kV}$, and the size of the supercell is $29.0292\,\text{\AA }\times 29.6278\,\text{\AA }\times 1000.678\,\text{\AA }$. The thickness of the sample is the same as in the case of relevant magnetic CBED simulations. The average value of the normalized TDS intensity at the location of magnetic reflections (red dashed circles) is $2.3\times 10^{-3}$, see Fig. 8.